OpenStudy (anonymous):

The siren of a fire engine that is driving northward at 30.0 m/s emits a sound of frequency 2000 Hz.A truck in front of this fire engine is moving northward at 20.0 m/s.What is the frequency of the siren's sound that the fire engine's driver hears reflected from the back of the truck?

OpenStudy (anonymous):

Hey Megan, doppler effect: apparent freq (n') of actual freq (n) n ' = n [v-vo] / [v-vs] ------------------------------------ in this formula, all v, (sound spped), vo (observer) and source (vs) are necesarily moving in the same direction (say + x). if one is not in + x, then is will be resolved therein. ------------------------------------- v=343 m/s (+x), vs = 30 m/s (+x), v0 = 20 m/s (+x) all in same direction the apparent freq to trucker n ' = 2000 [343 - 20] / [343 - 30] = 2063.90 Hz ----------------------------------- this n' is reflected and trucker becomes a virtual source vs = 20 (+x) , vo (engine) = 30 (+x) but sound is now directed from truck to engine, so resolved as v = - 343 (+x) n '' (apparent to engine) = n ' [- v - vo] / [- v - vs] n '' = n ' [v+vo] / [ v+vs] n '' = 2063.90 [343 +30] / [343 + 20] n '' = 2120.75 Hz this is what fire engine driver will hear as reflected freq Source: http://answers.yahoo.com/question/index?qid=20070503214006AA3oFMT Hope that helps, Good luck =)

OpenStudy (ujjwal):

The apparent frequency of sound due to doppler effect is given by: \[f_1=\frac{(relative. velocity. of. sound. and. listener)}{(relative. velocity. of. sound. and. source)}f \] where \(f\) is original frequency.